OpenStudy (anonymous):

Factor completely. -3k - k3 - 4k2

OpenStudy (lollylau):

those are powers right?

OpenStudy (anonymous):

i think

OpenStudy (lollylau):

factoring out -k, -k(3+k2+4k) =-k(k2+4k+3) =-k(k+1)(k+3)

OpenStudy (lollylau):

i'm bad at factoring D: but im fine with trinomials